# On Arratia’s coupling and the Dirichlet law for the factors of a random integer

### v2

@@ 8. Theorem 2 for the probabilistic model
Let $X_{1}>X_{2}>\cdots$ be the points of a Poisson process on $\mathbb{R}_{>0}$ with intensity $\frac{e^{-x}}{x}\,\mathrm{d}x$, and let $S\coloneqq\sum_{i\geqslant 1}X_{i}$. For each $i\geqslant 1$, Let $V_{i}\coloneqq\frac{X_{i}}{S}$. We then know that $(V_{1},V_{2},\ldots)$ follows the Poisson–Dirichlet distribution [9, Theorem 2.2]. With the Colouring Theorem (Proposition B.4), the point processes
@@
follows the distribution $\text{Dir}(\bm{\alpha})$ (see [11, Chapter 49, pp. 485-487] for a proof). ∎
9. A coupling for $\mathbf{D}_{x}$
Fix $k\geqslant 2$, $\bm{\alpha}\in\Delta^{k-1}$ with $\alpha_{i}>0$ for all $i$, and $f\in\mathcal{F}_{k}(\bm{\alpha})$. In addition, let $\mathbf{D}_{x}$ be a random variable satisfying (1.5), and let $\mathbf{Z}$ be a $\Delta^{k-1}$-valued random variable distributed according to $\text{Dir}(\bm{\alpha})$. We will construct an appropriate coupling between $\mathbf{D}_{x}$ and $\mathbf{Z}$. This will build on the coupling given in Section 2. To state our result, we also introduce the random vectors
(9.1) $\bm{\delta}_{x}\coloneqq\bigg{(}\frac{\log D_{x,1}}{\log N_{x}},\dots,\frac{\log D_{x,k}}{\log N_{x}}\bigg{)}\quad\text{and}\quad\bm{\delta}_{x}^{*}\coloneqq\left(\frac{\log D_{x,1}}{\log x},\ldots,\frac{\log D_{x,k}}{\log x}\right),$
which are more convenient to work with than $\mathbf{D}_{x}$ because we can compare them to $\mathbf{Z}$. To this end, we write $\left\|\cdot\right\|_{\infty}$ for the supremum norm in $\mathbb{R}^{k}$. Note that
(9.2) $\left\|\bm{\delta}_{x}-\bm{\delta}^{*}_{x}\right\|_{\infty}=\frac{\log(x/N_{x})}{(\log x)(\log N_{x})}\cdot\max_{1\leqslant i\leqslant k}\log D_{x,i}\leqslant\frac{\log(x/N_{x})}{\log x}.$
In addition, we have the following lemma:
@@ 9. A coupling for $\mathbf{D}_{x}$
Let $x\geqslant 2$ and assume the above notation. There exists a probability space $\Omega^{\prime}$ containing copies of the random variables $N_{x},\bm{\delta}_{x}$, $\bm{\delta}_{x}^{*}$, $\mathbf{V}$ and $\mathbf{Z}$, and an event $\mathcal{E}$ such that:
  1. (a)
    $\mathbb{P}[\mathcal{E}^{c}]\ll 1/\log x$;
  2. (b)
    If $\mathcal{E}$ occurs, then
    $\left\|\bm{\delta}_{x}-\mathbf{Z}\right\|_{\infty}\leqslant\frac{2\cdot\log(x/N_{x})+3\cdot\log s(N_{x})+2\cdot\Theta_{x}}{\log x}.$
@@
Our starting point is the coupling of $N_{x}$ and $\mathbf{V}$ described in Section 2. We equip this space with some additional random variables that are all independent of each other and of $N_{x}$ and $\mathbf{V}$, and whose role will become apparent later:
  • a sequence $(C_{i}^{\prime})_{i\geqslant 1}$ of random variables such that $\mathbb{P}[C_{i}^{\prime}=\ell]=\alpha_{\ell}$ for all $i\geqslant 1$ and all $\ell=1,2,\dots,k$;
  • for each natural number $n$ and each prime $p$, let $\big{(}X_{p}^{(1)}(n),\ldots,X_{p}^{(k)}(n)\big{)}$ be independent random vectors
    $\mathbb{P}\big{[}X_{p}^{(i)}(n)=e_{i}\ \forall i\leqslant k\big{]}=f(p^{e_{1}},\ldots,p^{e_{k}}),$
@@
Finally, we take =$\mathcal{E}=\{M_{x}=N_{x}\}$, which satisfies the required property (a) by Proposition 2.4.
We will now show that the space we constructed above contains copies of $\bm{\delta}_{x}$, $\bm{\delta}_{x}^{*}$ and of $\mathbf{Z}$ satisfying part (b). The construction of $\bm{\delta}_{x}$ and $\bm{\delta}_{x}^{*}$ is rather straightforward. Indeed, using the fact that $f\in\mathcal{F}_{k}(\bm{\alpha})$ (property (c) of Definition 1.1), we may easily check that the random $k$-tuple $(\prod_{p}p^{X_{p}^{(i)}(N_{x})})_{i=1}^{k}$ satisfies (1.5). Thus, we may take
$D_{x,i}\coloneqq\prod_{p}p^{X_{p}^{(i)}(N_{x})}\quad\text{for}\ i=1,2,\dots,k,$
so that the distribution of $\mathbf{D}_{x}=(D_{x,1},\dots,D_{x,k})$ is indeed in accordance with (1.5). We then define $\bm{\delta}_{x}$ and $\bm{\delta}_{x}^{*}$ as in (9.1).
Next, we define $\mathbf{Z}$. Firstly, we introduce a new sequence of random variables $(C_{i})_{i\geqslant 1}$. To define them, fix $n\in\mathbb{Z}\cap[1,x]$ and, for every prime $p|n^{\flat}=n/s(n)$, fix a number $\ell(p)\in\{1,\dots,k\}$. Moreover, fix choices of $\ell_{1}^{\prime},\ell_{2}^{\prime},\ldots\in\{1,2,\dots,k\}$. Recall that $N_{x}=P_{1}P_{2}\cdots$ with the $P_{i}$’s forming a non-increasing sequence of primes or ones. Assume that the events $N_{x}=n$, $X_{p}^{(j)}=\mathbbm{1}_{j=\ell(p)}$ for $j=1,\dots,k$, and $C_{i}^{\prime}=\ell_{i}^{\prime}$ occur.
  • If $i\geqslant 1$ is such that $P_{i}|n^{\flat}$, then we let $C_{i}=\ell(P_{i})$.
  • If $i\geqslant 1$ is such that $P_{i}|s(n)$, then we let $C_{i}=\ell_{i}^{\prime}$.
Since $f\in\mathcal{F}_{k}(\bm{\alpha})$ satisfies property (b) of Definition 1.1, a straightforward computation reveals that the random variables $(C_{i})_{i\geqslant 1}$ are independent of each other, of $N_{x}$ and of $\mathbf{V}$, and they satisfy
@@
$\mathbf{Z}\coloneqq\left(\sum_{i\geqslant 1:\ C_{i}=1}V_{1},\ldots,\sum_{i\geqslant 1:\ C_{i}=k}V_{k}\right),$
@@
In order to complete the proof of the lemma, it remains to show that (b) holds when =$\mathcal{E}=\{M_{x}=N_{x}\}$. Define the auxiliary random variable
$\bm{\rho}_{x}\coloneqq\left(\sum_{i\geqslant 1:\ C_{i}=1}\frac{\log P_{i}}{\log x},\ldots,\sum_{i\geqslant 1:\ C_{i}=k}\frac{\log P_{i}}{\log x}\right).$
When $\mathcal{E}$ holds, then Lemma 2.1 implies that
Thus, we readily find that
(9.3) $\left\|\bm{\rho}_{x}-\mathbf{Z}\right\|_{\infty}\leqslant\frac{\log(x/N_{x})+2\cdot\log s(N_{x})+2\cdot\Theta_{x}}{\log x}.$
In addition, by the construction of the random variables $C_{i}$, we claim that
(9.4) $\left\|\bm{\delta}_{x}^{*}-\bm{\rho}_{x}\right\|_{\infty}\leqslant\frac{\log s(N_{x})}{\log x}.$
Indeed, if $N_{x}=n$, then we have
$\log D_{x,\ell}$ $=\sum_{p|n^{\flat}}X_{p}^{(\ell)}(n)\log p+\sum_{p|s(n)}X_{p}^{(\ell)}(n)\log p$
$=\sum_{i\geqslant 1:\ P_{i}|n^{\flat}}\mathbbm{1}_{X_{P_{i}}^{(\ell)}(n)=1}\log P_{i}+\sum_{p|s(n)}X_{p}^{(\ell)}(n)\log p$
$=\sum_{i\geqslant 1:\ P_{i}|n^{\flat}}\mathbbm{1}_{C_{i}=\ell}\log P_{i}+\sum_{p|s(n)}X_{p}^{(\ell)}(n)\log p,$
@@
Combining (9.2)-(9.4) completes the proof of part (b), and hence of the lemma. ∎
@@
We write $B(\mathbf{x},r)$ for the closed ball of radius $r\geqslant 0$ centered at $\mathbf{x}$ in the normed vector space $(\mathbb{R}^{k},\left\|\cdot\right\|_{\infty})$. We recall that $\Delta^{k-1}$ is the standard $(k-1)$-dimensional simplex. For any subset $A\subseteq\Delta^{k-1}$, we write $\partial A$ for the boundary of $A$ in the relative topology of $\Delta^{k-1}$ (cf. Definition C.1(c)).
@@
  • $\mathcal{E}_{\mathrm{NT}}(A)\coloneqq\big{\{}B(\bm{\delta}^{*}_{x},R_{\mathrm{NT}})\cap\partial A=\emptyset\big{\}}$;
@@
$\mathbb{P}\big{[}\bm{\delta}_{x}\in A\big{]}=\mathbb{P}\big{[}\mathbf{Z}\in A\big{]}+O\left(\mathbb{P}\big{[}\mathcal{E}_{\mathrm{NT}}(A)^{c}\big{]}+\mathbb{P}\big{[}\mathcal{E}_{\mathrm{PD}}(A)^{c}\big{]}+\frac{1}{\log x}\right).$
@@
(9.5) |[$\Big{|}\mathbb{P}\big{[}\mathcal{E}\cap\{\bm{\delta}_{x}\in A\}\big{]}-\mathbb{P}\big{[}\mathcal{E}\cap\{\mathbf{Z}\in A\}\big{]}\Big{|}\leqslant\mathbb{P}\big{[}\mathcal{E}_{\mathrm{NT}}(A)^{c}\big{]}+\mathbb{P}\big{[}\mathcal{E}_{\mathrm{PD}}(A)^{c}\big{]}.$
We shall prove the following stronger statement: if $\mathcal{E}\cap\mathcal{E}_{\mathrm{NT}}(A)\cap\mathcal{E}_{\mathrm{PD}}(A)$ occurs, then $\bm{\delta}_{x}\in A$ if, and only if, $\mathbf{Z}\in A$.
Indeed, assume that $\mathcal{E}\cap\mathcal{E}_{\mathrm{NT}}(A)\cap\mathcal{E}_{\mathrm{PD}}(A)$ occurs, and we also have $\bm{\delta}_{x}\in A$. Aiming for a contradiction, we also assume that $\mathbf{Z}\notin A$.
Let $R_{\mathrm{NT}}^{\prime}\coloneqq(\log x)^{-1}(2\cdot\log(x/N_{x})+3\cdot\log s(N_{x}))$. Note that $B(\bm{\delta}_{x},R_{\mathrm{NT}}^{\prime})\subseteq B(\bm{\delta}_{x}^{*},R_{\mathrm{NT}})$ by (9.2). By assumption of $\mathcal{E}_{\mathrm{NT}}(A)$, we know that $B(\bm{\delta}_{x},R_{\mathrm{NT}}^{\prime})\cap\partial A=\emptyset$. Since $\bm{\delta}_{x}\in A$, we use Lemma C.2 to find
$B(\bm{\delta}_{x},R_{\mathrm{NT}}^{\prime})\cap\Delta^{k-1}\subseteq A.$
@@
In particular, we find that $B(\bm{\delta}_{x},R_{\mathrm{NT}}^{\prime})\cap B(\mathbf{Z},R_{\mathrm{PD}})\cap\Delta^{k-1}=\emptyset$. On the other hand, we know that $\left\|\bm{\delta}_{x}-\mathbf{Z}\right\|_{\infty}\leqslant R_{\mathrm{NT}}^{\prime}+R_{\mathrm{PD}}$ when $\mathcal{E}$ occurs by Lemma 9.1(b). In particular, there exists a point $\mathbf{y}$ on the line segment connecting $\bm{\delta}_{x}$ and $\mathbf{Z}$ such that $\left\|\bm{\delta}_{x}-\mathbf{y}\right\|_{\infty}\leqslant R_{\mathrm{NT}}^{\prime}$ and $\left\|\mathbf{y}-\mathbf{Z}\right\|_{\infty}\leqslant R_{\mathrm{PD}}$. Since the sets $\Delta^{k-1}$, $B(\bm{\delta}_{x},R_{\mathrm{NT}}^{\prime})$ and $B(\mathbf{Z},R_{\mathrm{PD}})$ are convex, we conclude that $\mathbf{y}$ lies in their intersection. But we had seen before that this intersection is the empty set. We have thus arrived at a contradiction. This completes the proof that if $\mathcal{E}\cap\mathcal{E}_{\mathrm{NT}}(A)\cap\mathcal{E}_{\mathrm{PD}}(A)$ occurs, and we also know that $\bm{\delta}_{x}\in A$, then we must also have that $\mathbf{Z}\in A$.
Conversely, we may show by a simple variation of the above argument that if $\mathcal{E}\cap\mathcal{E}_{\mathrm{NT}}(A)\cap\mathcal{E}_{\mathrm{PD}}(A)$ occurs, and we also know that $\mathbf{Z}\in A$, then we must also have that $\bm{\delta}_{x}\in A$. This completes the proof of the lemma. ∎
@@
$\mathcal{E}_{\mathrm{NT}}(\Delta^{k-1}_{\mathbf{u}})^{c}\subseteq\bigcup_{\begin{subarray}{c}i<k\\ u_{i}\neq 1\end{subarray}}\mathcal{B}_{\mathrm{NT}}^{(i)}(u_{i}),\quad\text{where}\quad\mathcal{B}_{\mathrm{NT}}^{(i)}(u)\coloneqq\Big{\{}\big{|}\log D_{x,i}-u\log x\big{|}\leqslant 3\log(x/N_{x}^{\flat})\Big{\}},$
@@
(9.6) $\mathbb{P}[\mathbf{D}_{x}\in A]=\mathbb{P}[\mathbf{Z}\in A]+O\left(\mathbb{P}\big{[}\mathcal{E}^{*}(A)^{c}\big{]}+\frac{1}{\log x}\right)$
for any measurable set $A$. With Chernoff’s bound and Lemmas 2.2-2.3, the probability that $R^{*}>\log_{2}x$ is $\ll 1/\log x$. Moreover, for any $\delta\in(0,1/4]$, we can show by a direct computation with the Dirichlet distribution the uniform bound
@@
$\mathbb{P}\big{[}D_{x,i}\leqslant N_{x}^{u_{i}}\ \,\forall i<k\big{]}=F_{\bm{\alpha}}(\mathbf{u})+O\Bigg{(}\sum_{\begin{subarray}{c}1\leqslant i<k\\ u_{i}\neq 1\end{subarray}}\frac{1}{(1+u_{i}\frac{\log x}{\log_{2}x})^{1-\alpha_{i}}(1+(1-u_{i})\frac{\log x}{\log_{2}x})^{\alpha_{i}}}\Bigg{)}.$
@@
Relation (1.4) implies that $f(d_{1}^{\prime},\dots,d_{k}^{\prime})=\prod_{j=1}^{k}\alpha_{j}^{\omega(d_{j}^{\prime})}$. Moreover, using property (a) of Definition 1.1, we find that the sum over $d_{j}^{\prime\prime}$ equals $1$. In conclusion,
@@
Because this paper lies in the intersection of number theory and probability theory, readers coming from one of these fields might not be familiar with standard results of the other one. For this reason, we gather here some key results from both fields. We shall also need a standard fact from topology. We present these results in the following three sections.

### v3

@@ 8. Theorem 2 for the probabilistic model
Let $X_{1}>X_{2}>\cdots$ be the points of a Poisson process on $\mathbb{R}_{>0}$ with intensity $\frac{e^{-x}}{x}\,\mathrm{d}x$, and let $S\coloneqq\sum_{i\geqslant 1}X_{i}$. For each $i\geqslant 1$, Let $V_{i}\coloneqq\frac{X_{i}}{S}$. We then know that $(V_{1},V_{2},\ldots)$ follows the Poisson–Dirichlet distribution [11, Theorem 2.2]. With the Colouring Theorem (Proposition B.4), the point processes
@@
follows the distribution $\text{Dir}(\bm{\alpha})$ (see [13, Chapter 49, pp. 485-487] for a proof). ∎
9. A coupling for $\mathbf{D}_{f,x}$
Fix $k\geqslant 2$, $\bm{\alpha}\in\Delta^{k-1}$ with $\alpha_{i}>0$ for all $i$, and $f\in\mathcal{F}_{k}(\bm{\alpha})$. In addition, let $N_{x}$ be a random integer uniformly distributed in $\mathbb{Z}\cap[1,x]$, let $\mathbf{D}_{f,x}$ be a random vector satisfying (1.7), and let $\mathbf{Z}$ be a $\Delta^{k-1}$-valued random variable distributed according to $\text{Dir}(\bm{\alpha})$. We also consider the random vector $\bm{\delta}_{f,x}\in\Delta^{k-1}$ defined by
(9.1) $\bm{\delta}_{f,x}\coloneqq\bigg{(}\frac{\log D_{f,x,1}}{\log N_{x}},\dots,\frac{\log D_{f,x,k}}{\log N_{x}}\bigg{)}$
when $N_{x}\geqslant 2$, and $\bm{\delta}_{f,x}=(1,0,\ldots,0)$ when $N_{x}=1$. We note that Theorem 2 is about comparing the cumulative distribution function of $\bm{\delta}_{f,x}$ with the one for $\mathbf{Z}$. To achieve this, we construct an appropriate coupling between $N_{x}$, $\mathbf{D}_{f,x}$ and $\mathbf{Z}$. This will build on the coupling given in Section 2.
@@ 9. A coupling for $\mathbf{D}_{f,x}$
Let $x\geqslant 2$ and assume the above notation. There exists a coupling of $N_{x}$, $\mathbf{V}$, $\mathbf{D}_{f,x}$ and $\mathbf{Z}$ such that $\mathbb{P}[\mathcal{E}^{c}]\ll\frac{1}{\log x}$ with
$\mathcal{E}\coloneqq\Bigg{\{}\left\|\bm{\delta}_{f,x}-\mathbf{Z}\right\|_{\infty}\leqslant\frac{2\cdot\log(x/N_{x})+3\cdot\log s(N_{x})+2\cdot\Theta_{x}}{\log x}\Bigg{\}},$
with $\bm{\delta}_{f,x}$ defined as in (9.1), $\Theta_{x}$ defined as in (2.5), and $\left\|\cdot\right\|_{\infty}$ being the supremum norm in $\mathbb{R}^{k}$.
@@
Our starting point is the coupling of $N_{x}$ and $\mathbf{V}$ described in Section 2, and it remains to construct $\mathbf{D}_{f,x}$ and $\mathbf{Z}$. To do this, we need to equip this space with the following additional random variables that are all independent of each other and of $N_{x}$ and $\mathbf{V}$, and whose role will become apparent later:
  • a sequence $(C_{i}^{\prime})_{i\geqslant 1}$ of i.i.d. random variables such that $\mathbb{P}[C_{i}^{\prime}=\ell]=\alpha_{\ell}$ for all $i\geqslant 1$ and all $\ell=1,2,\dots,k$;
  • for each natural number $n$ and each prime $p$, let $\mathbf{X}_{p}(n)\coloneqq\big{(}X_{p,1}(n),\ldots,X_{p,k}(n)\big{)}$ be independent random vectors with
    $\mathbb{P}\big{[}X_{p,i}(n)=e_{i}\ \forall i\leqslant k\big{]}=f(p^{e_{1}},\ldots,p^{e_{k}}),$
@@
To construct $\mathbf{D}_{f,x}$, we may take
$D_{f,x,i}\coloneqq\prod_{p}p^{X_{p,i}(N_{x})}\quad\text{for}\ i=1,2,\dots,k.$
Since $f\in\mathcal{F}_{k}(\bm{\alpha})$ satisfies property (c) of Definition 1.2, we may easily check that the distribution of $\mathbf{D}_{f,x}=(D_{f,x,1},\dots,D_{f,x,k})$ is indeed in accordance with (1.7).
Next, we construct $\mathbf{Z}$. Firstly, we need to construct a new sequence of random variables $(C_{i})_{i\geqslant 1}$, whose definition depends deterministically on the previously introduced random variables. The following events must occur:
  • We have $N_{x}=n$ for some choice of $n\in\mathbb{Z}\cap[1,x]$;
  • For all $j\in\{1,\ldots,k\}$ and all $p|n^{\flat}$, we have $X_{p,j}(n)=\mathbbm{1}_{j=\ell(p)}$ for some choice of sequence $(\ell(p))_{p|n^{\flat}}$ taking values in $\{1,\ldots,k\}$.
  • For all $i\geqslant 1$, we have that $C^{\prime}_{i}=\ell^{\prime}_{i}$ for some choice sequence $(\ell^{\prime}_{i})_{i\geqslant 1}$ taking values in $\{1,\ldots,k\}$.
Recall the factorization $N_{x}=P_{1}P_{2}\cdots$ with the $P_{i}$’s forming a non-increasing sequence of primes or ones. Then we construct $(C_{i})_{i\geqslant 1}$ in the following way:
  • If $i\geqslant 1$ is such that $P_{i}>1$ and $P_{i}|n^{\flat}$, then we let $C_{i}\coloneqq\ell(P_{i})$.
  • Otherwise, we let $C_{i}\coloneqq\ell_{i}^{\prime}$.
Since $f\in\mathcal{F}_{k}(\bm{\alpha})$ satisfies property (b) of Definition 1.2, a straightforward computation reveals that for any $n\in\mathbb{Z}\cap[1,x]$, any $(\ell_{i})_{i\in I}$ taking values in $\{1,\ldots,k\}$ and any finite set $I\subset\mathbb{N}$, we have
$\mathbb{P}\big{[}C_{i}=\ell_{i}\ \ \forall i\in I\,|\,N_{x}=n\big{]}=\prod_{j=1}^{k}\alpha_{j}^{\#\{i\in I:\ell_{i}=j\}},$
implying that the random variables $(C_{i})_{i\geqslant 1}$ are independent of each other, of $N_{x}$ and of $\mathbf{V}$, and they satisfy
@@
$\mathbf{Z}\coloneqq\left(\sum_{i:\,C_{i}=1}V_{1},\ldots,\sum_{i:\,C_{i}=k}V_{k}\right),$
@@
Recall the definition of $\bm{\delta}_{f,x}$ in (9.1). We introduce the auxiliary random variables
(9.2) $\bm{\delta}_{f,x}^{*}\coloneqq\left(\frac{\log D_{f,x,1}}{\log x},\ldots,\frac{\log D_{f,x,k}}{\log x}\right),\qquad\bm{\rho}_{f,x}\coloneqq\left(\sum_{i:\,C_{i}=1}\frac{\log P_{i}}{\log x},\ldots,\sum_{i:\,C_{i}=k}\frac{\log P_{i}}{\log x}\right).$
We claim that
(9.3) $\left\|\bm{\delta}_{f,x}-\bm{\delta}_{f,x}^{*}\right\|_{\infty}\leqslant\frac{\log(x/N_{x})}{\log x}.$
Indeed, if $N_{x}=1$, both the left-hand and right-hand sides of the inequality are always equal to 1. For $N_{x}\geqslant 2$, we note that
$\left\|\bm{\delta}_{f,x}-\bm{\delta}_{f,x}^{*}\right\|_{\infty}=\frac{\log(x/N_{x})}{\log x}\cdot\max_{1\leqslant i\leqslant k}\frac{\log D_{f,x,i}}{\log N_{x}},$
from which (9.3) follows.
Furthermore, we also assert that
(9.4) $\left\|\bm{\delta}_{f,x}^{*}-\bm{\rho}_{f,x}\right\|_{\infty}\leqslant\frac{\log s(N_{x})}{\log x}.$
To verify this, suppose $N_{x}=n$. Then we have
$\log D_{x,\ell}$ $=\sum_{p|n^{\flat}}X_{p,\ell}(n)\log p+\sum_{p|s(n)}X_{p,\ell}(n)\log p$
$=\sum_{i\geqslant 1:\ P_{i}|n^{\flat}}X_{P_{i},\ell}(n)\log P_{i}+\sum_{p|s(n)}X_{p,\ell}(n)\log p$
$=\sum_{i\geqslant 1:\ P_{i}|n^{\flat}}\mathbbm{1}_{C_{i}=\ell}\log P_{i}+\sum_{p|s(n)}X_{p,\ell}(n)\log p,$
@@
Now if the event $\mathcal{E}$ does not happen, then
$\left\|\bm{\delta}_{f,x}-\mathbf{Z}\right\|_{\infty}>\frac{2\cdot\log(x/N_{x})+3\cdot\log s(N_{x})+2\cdot\Theta_{x}}{\log x}.$
Together with (9.3) and (9.4), this implies that
$\left\|\bm{\rho}_{f,x}-\mathbf{Z}\right\|_{\infty}>\frac{\log(x/N_{x})+2\cdot\log s(N_{x})+2\cdot\Theta_{x}}{\log x}$
By Lemma 2.1, we must then have that $M_{x}\neq N_{x}$. In conclusion, we have proven that $\mathcal{E}^{c}\subseteq\{M_{x}\neq N_{x}\}$. Thus, Proposition 2.4 shows that $\mathbb{P}[\mathcal{E}^{c}]\ll\frac{1}{\log x}$, as required. ∎
@@
We write $B(\mathbf{x},r)$ for the closed ball of radius $r\geqslant 0$ centered at $\mathbf{x}$ in the normed vector space $(\mathbb{R}^{k},\left\|\cdot\right\|_{\infty})$. We recall that $\Delta^{k-1}$ is the standard $(k-1)$-dimensional simplex. For any subset $A\subseteq\Delta^{k-1}$, we write $\partial A$ for the boundary of $A$ in the relative topology of $\Delta^{k-1}$.
@@
  • $\mathcal{E}_{\mathrm{NT}}(A)\coloneqq\big{\{}B(\bm{\delta}_{f,x}^{*},R_{\mathrm{NT}})\cap\partial A=\emptyset\big{\}}$;
@@
$\mathbb{P}\big{[}\bm{\delta}_{f,x}\in A\big{]}=\mathbb{P}\big{[}\mathbf{Z}\in A\big{]}+O\left(\mathbb{P}\big{[}\mathcal{E}_{\mathrm{NT}}(A)^{c}\big{]}+\mathbb{P}\big{[}\mathcal{E}_{\mathrm{PD}}(A)^{c}\big{]}+\frac{1}{\log x}\right).$
@@
(9.5) |[$\Big{|}\mathbb{P}\big{[}\mathcal{E}\cap\{\bm{\delta}_{f,x}\in A\}\big{]}-\mathbb{P}\big{[}\mathcal{E}\cap\{\mathbf{Z}\in A\}\big{]}\Big{|}\leqslant\mathbb{P}\big{[}\mathcal{E}_{\mathrm{NT}}(A)^{c}\big{]}+\mathbb{P}\big{[}\mathcal{E}_{\mathrm{PD}}(A)^{c}\big{]}.$
We shall prove the following stronger statement: if $\mathcal{E}\cap\mathcal{E}_{\mathrm{NT}}(A)\cap\mathcal{E}_{\mathrm{PD}}(A)$ occurs, then $\bm{\delta}_{f,x}\in A$ if, and only if, $\mathbf{Z}\in A$.
Indeed, assume that $\mathcal{E}\cap\mathcal{E}_{\mathrm{NT}}(A)\cap\mathcal{E}_{\mathrm{PD}}(A)$ occurs, and we also have $\bm{\delta}_{f,x}\in A$. Aiming for a contradiction, we also assume that $\mathbf{Z}\notin A$.
Let $R_{\mathrm{NT}}^{\prime}\coloneqq(\log x)^{-1}(2\cdot\log(x/N_{x})+3\cdot\log s(N_{x}))$. Note that $B(\bm{\delta}_{f,x},R_{\mathrm{NT}}^{\prime})\subseteq B(\bm{\delta}_{f,x}^{*},R_{\mathrm{NT}})$ by (9.3). By assumption of $\mathcal{E}_{\mathrm{NT}}(A)$, we know that $B(\bm{\delta}_{f,x},R_{\mathrm{NT}}^{\prime})\cap\partial A=\emptyset$. Since $\bm{\delta}_{f,x}\in A$ and $B(\bm{\delta}_{f,x},R_{\mathrm{NT}}^{\prime})\cap\Delta^{k-1}$ is connected (being the intersection of two convex sets), we have that
$B(\bm{\delta}_{f,x},R_{\mathrm{NT}}^{\prime})\cap\Delta^{k-1}\subseteq A.$
@@
In particular, we find that $B(\bm{\delta}_{f,x},R_{\mathrm{NT}}^{\prime})\cap B(\mathbf{Z},R_{\mathrm{PD}})\cap\Delta^{k-1}=\emptyset$. On the other hand, we know that $\left\|\bm{\delta}_{f,x}-\mathbf{Z}\right\|_{\infty}\leqslant R_{\mathrm{NT}}^{\prime}+R_{\mathrm{PD}}$ when $\mathcal{E}$ occurs by Lemma 9.1(b). In particular, there exists a point $\mathbf{y}$ on the line segment connecting $\bm{\delta}_{f,x}$ and $\mathbf{Z}$ such that $\left\|\bm{\delta}_{f,x}-\mathbf{y}\right\|_{\infty}\leqslant R_{\mathrm{NT}}^{\prime}$ and $\left\|\mathbf{y}-\mathbf{Z}\right\|_{\infty}\leqslant R_{\mathrm{PD}}$. Since the sets $\Delta^{k-1}$, $B(\bm{\delta}_{f,x},R_{\mathrm{NT}}^{\prime})$ and $B(\mathbf{Z},R_{\mathrm{PD}})$ are convex, we conclude that $\mathbf{y}$ lies in their intersection. But we had seen before that this intersection is the empty set. We have thus arrived at a contradiction. This completes the proof that if $\mathcal{E}\cap\mathcal{E}_{\mathrm{NT}}(A)\cap\mathcal{E}_{\mathrm{PD}}(A)$ occurs, and we also know that $\bm{\delta}_{f,x}\in A$, then we must also have that $\mathbf{Z}\in A$.
Conversely, we may show by a simple variation of the above argument that if $\mathcal{E}\cap\mathcal{E}_{\mathrm{NT}}(A)\cap\mathcal{E}_{\mathrm{PD}}(A)$ occurs, and we also know that $\mathbf{Z}\in A$, then we must also have that $\bm{\delta}_{f,x}\in A$. This completes the proof of the lemma. ∎
@@
$\mathcal{E}_{\mathrm{NT}}(\Delta^{k-1}_{\mathbf{u}})^{c}\subseteq\bigcup_{\begin{subarray}{c}i<k\\ u_{i}\neq 1\end{subarray}}\mathcal{B}_{\mathrm{NT}}^{(i)}(u_{i}),\quad\text{where}\quad\mathcal{B}_{\mathrm{NT}}^{(i)}(u)\coloneqq\Big{\{}\big{|}\log D_{f,x,i}-u\log x\big{|}\leqslant 3\log(x/N_{x}^{\flat})\Big{\}},$
@@
(9.6) $\mathbb{P}[\mathbf{D}_{f,x}\in A]=\mathbb{P}[\mathbf{Z}\in A]+O\left(\mathbb{P}\big{[}\mathcal{E}^{*}(A)^{c}\big{]}+\frac{1}{\log x}\right)$
for any measurable set $A$. With Chernoff’s bound and Lemmas 2.2-2.3, the probability that $R^{*}>\frac{\log_{2}x}{\log x}$ is $\ll 1/\log x$. Moreover, for any $\delta\in(0,1/4]$, we can show by a direct computation with the Dirichlet distribution the uniform bound
@@
$\mathbb{P}\big{[}D_{f,x,i}\leqslant N_{x}^{u_{i}}\ \,\forall i<k\big{]}=F_{\bm{\alpha}}(\mathbf{u})+O\Bigg{(}\sum_{\begin{subarray}{c}1\leqslant i<k\\ u_{i}\neq 1\end{subarray}}\frac{1}{(1+u_{i}\frac{\log x}{\log_{2}x})^{1-\alpha_{i}}(1+(1-u_{i})\frac{\log x}{\log_{2}x})^{\alpha_{i}}}\Bigg{)}.$
@@
Relation (1.6) implies that $f(d_{1}^{\prime},\dots,d_{k}^{\prime})=\prod_{j=1}^{k}\alpha_{j}^{\omega(d_{j}^{\prime})}$. Moreover, using property (a) of Definition 1.2, we find that the sum over the $d_{j}^{\prime\prime}$’s equals $1$. In conclusion,
@@
Because this paper lies in the intersection of number theory and probability theory, readers coming from one of these fields might not be familiar with standard results of the other one. For this reason, we gather here some key results from both fields. We present these results in the following two sections.